m=2,(6,5) Find a Equation
Answer by lefty4ever26(59)

(y-y,)=m(x-x,) You are given m and the point to plug into the formula.
So you get y-5=2(x-6) Simplify
y-5=2x-12
y=2x-7 this is the equation.